### v1

# Flipping Out: Role of Arginine in Hydrophobic Polymer Collapse

Abstract
Arginine has been a mainstay in biological formulation development for decades. To date, the way arginine modulates protein stability has been widely studied and debated. Here, we employed a hydrophobic polymer to decouple hydrophobic effects from other interactions relevant to protein folding. While existing hypotheses for the effects of arginine can generally be categorized as either direct or indirect, our results indicate that direct and indirect mechanisms of arginine co-exist and oppose each other. At low concentrations, arginine was observed to stabilize hydrophobic polymer collapse via a sidechain-dominated direct mechanism, while at high concentrations, arginine stabilized polymer collapse via a backbone-dominated indirect mechanism. These findings highlight the modular nature of the widely used additive arginine, with relevance in the design of stable biological formulations.
keywords:
Hydrophobic interactions, formulation design, excipient, osmolyte, protein stability, protein folding
[…]
{tocentry}
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification. 7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation 10, 11, 12, denaturation13, 14, 15, and activation16 of certain proteins and viruses. Additional studies have found that arginine increases or decreases protein melting temperature depending on concentration 17, 18. Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
The molecular mechanism for arginine effects on protein stability remains elusive. In general, additive effects on protein stability are thought to be either direct or indirect. Direct mechanisms involve direct protein-additive interactions that drive stabilization, while indirect mechanisms influence protein stability by modulating the surrounding solvent structure. It remains debated whether arginine acts primarily via a direct or indirect mechanism. Several studies called attention to direct interactions between arginine and aromatic residues 7, 19, 20, acidic residues 12, 17, and hydrophobic moieties 21, 22. Other studies have proposed clusters of free arginine molecules in solution enable the crowding out of protein-protein interactions 23, 20, 24, 17, or alteration of hydration shell water dynamics.25, 26 The wide range of observations related to the role of arginine on protein stability suggests that arginine harbors diverse, context-dependent mechanisms.
To elucidate the mechanisms through which arginine influences protein stability, this study focuses on its effects on hydrophobic interactions. Hydrophobic interactions are key in several biologically relevant phenomena, including protein folding and stability.27, 28, 29, 30, 31, 32, 33, 34, 35 Additives in solutions have been observed to modulate the strength of hydrophobic interactions, and in turn, the stability of proteins.36, 37, 38, 39, 40, 41, 42 For example, simulation studies have shown that trimethylamine N-oxide (TMAO) has a negligible effect on or strengthens hydrophobic interactions,38, 43, 44, 45, 46, 47 while hydrophobic interactions are weakened in urea solutions.48, 49, 37, 50, 51, 39, 40, 52 Indeed, the effects of urea and TMAO on hydrophobic interactions are consistent with their experimentally-observed roles as a protein denaturant and stabilizer, respectively.53, 54, 55, 41
Molecular dynamics (MD) simulations have provided valuable insights towards understanding these mechanisms as they relate to hydrophobicity. Several studies have highlighted the utilization of a hydrophobic polymer model for describing the role of solvent and additives on protein-like collapse. 56, 40, 57, 58, 39, 59 The use of a hydrophobic polymer model enables the decoupling of additive effects on hydrophobic vs other interactions, which is challenging in experiments. In the present study, we utilize MD simulations of a hydrophobic polymer to characterize the effects of arginine on many-body hydrophobic interactions pertinent to protein stability.
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). Replica exchange umbrella sampling60 (REUS) simulations were performed using GROMACS 2021.461, 62 with the PLUMED 2.8.0 63, 64 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$38. Chloride (${}^{-}$) counterions were added to achieve charge neutrality. The TIP4P/200565 model was used to describe water, and the CHARM22 force field was used for arginine and ${}^{-}$ 66. Lorentz-Berthelot mixing rules67 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
The potential of mean force (PMF) along the radius of gyration of the polymer was calculated as ${W(R_{g})=-k_{B}Tln(P(R_{g}))}$. Biased probability distributions were reweighted according to the Weighted Histogram Analysis Method (WHAM).68 The free energy of polymer unfolding ($\mathrm{\Delta G_{u}}$) was calculated according to:
[…]
where $\mathrm{R_{g,cut}}$ was determined as the point between the folded and unfolded states where $\frac{\partial W(R_{g})}{\partial R_{g}}=0$.
Fig. 1a shows the PMF along the $R_{g}$ reaction coordinate. In all solutions, free energy minima were observed at approximately 0.4 and 0.8 nm (labeled as III in Fig. 1g), along with a prominent free energy barrier at $\sim$0.6 nm representing the transition between folded and unfolded states. (Fig. 1a). In pure water, hydrophobic collapse is unfavorable, with the unfolded state favored by $\sim$0.3 kT. In contrast, at all arginine concentrations, the folded state of the polymer is favored relative to pure water, and a monotonic increase in $\Delta G_{u}$ is observed (Fig. S4). An additional barrier at $\sim$0.45 nm was identified separating two folded states, labeled as I and II in Fig. 1a.
The folded polymer ensemble in arginine solutions exhibits free energy minima corresponding to globular ($\sim$0.4 nm) and hairpin-like ($\sim$0.5 nm) configurations (Fig. 1g). Arginine clusters encapsulating the hydrophobic polymer are observed in each state (Fig. S6). We propose that the free energy barrier separating these two states arises from an energetic penalty associated with breaking these encapsulating clusters. Such a mechanism is similar to that observed by Li et al.22, who observed arginine-mediated suppression of hydrophobic association.
The PMF was decomposed into individual components to further investigate the role of arginine in polymer collapse. Following the methods outlined by several others,69, 70, 42, 71 the PMF was decomposed as
$\begin{split}W(R_{g})&=W_{vac}(R_{g})+W_{cav}(R_{g})\\ &\quad+E_{pw}(R_{g})+E_{pa}(R_{g})\\ &\quad+E_{pc}(R_{g})\end{split}$ (2)
$W_{vac}(R_{g})$ captures intrapolymer degrees of freedom and was obtained from independent REUS simulations of the polymer in vacuum. $E_{pw}(R_{g})$, $E_{pa}(R_{g})$, and $E_{pc}(R_{g})$ are average polymer-water, polymer-arginine, and polymer-chloride interaction energies, respectively. The remaining term is $W_{cav}(R_{g})$, which is the cavitation component and quantifies the energetic cost of forming a cavity – of the same size and shape as the polymer – in the solution.
Figure 1: PMF decomposition in 0.0 M, 0.25 M, 0.5 M, and 1.0 M arginine solutions. (a) The PMF obtained along the $R_{g}$ reaction coordinate, $W(R_{g})$, (b) vacuum contribution, $W_{vac}$, (c) cavitation contribution, $W_{cav}$, (d) polymer-water interactions, $E_{pw}$, (e) polymer-arginine interactions, $E_{pa}$, and (f) polymer-chloride interactions, $E_{pc}$. (g) Representative configurations along the reaction coordinate as denoted in (a) as I, II, and III. (h) Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), cavitation contribution ($\Delta\Delta G_{cav}$), polymer-water interactions ($\Delta\Delta E_{pw}$), and polymer-arginine interactions ($\Delta\Delta E_{pa}$). Increasing arginine concentration is denoted by increased shading (light to dark) and is indicated by arrows in (a-f). The polymer in water alone is shown in black, where appropriate. Mean values were estimated from three replicate REUS simulations. Error bars are reported as described in the SI. All plots are normalized to 0 at $R_{g}=0.4$ nm, where appropriate.
Fig. 1b-f shows the decomposition of the PMF into these various components. The vacuum component favors the folded state of the polymer, associated with favorable intrapolymer interactions and configurational entropy upon collapse (Fig. 1b). Because this component does not depend on the presence of arginine, a balance of the remaining components dictates the effect of arginine on hydrophobic polymer collapse.
Near large idealized solutes or hydrophobic interfaces, water dewets and forms a vapor-liquid-like interface72, 73. Similar behavior may be expected for our hydrophobic polymer; hence, we computed $W_{cav}$, which depends on both the size and shape of the polymer and is related to the air-water surface tension70, 74, 69. The cavitation component favors the folded state (Fig. 1c), reflecting a strong hydrophobic driving force for polymer collapse56.
Attractive polymer-water interactions become more favorable with increasing $R_{g}$ (Fig. 1d), indicating polymer-water attractive interactions oppose polymer collapse. It is worth noting that the free energy minima at $\sim$0.8 nm in the unfolded ensemble is observed in an aqueous environment (Fig. 1a), but not in vacuum (Fig. 1b). This minima arises due to favorable polymer-water interactions, consistent with prior MD simulations that observed water-mediated interactions drive large hydrophobic solutes apart.75, 76, 57. Additionally, sufficient dewetting of the hydrophobic polymer is a known bottleneck to collapse,77, 56, 69 resulting in the free energy barrier at $\sim$0.6 nm separating the folded and unfolded states.
Attractive polymer-arginine interactions approach an energetic minima at $\sim$0.5 nm (Fig. 1e), giving rise to the global minimum observed in the overall PMF. Polymer-chloride interactions were observed on the order of thermal fluctuations ($<0.5$ kT), consistent with previous observations that chloride ions are depleted from the local domain of hydrophobic solutes.78, 37
Fig. 1h shows the change in each component upon unfolding in arginine solution relative to that observed in water. In Fig. 1h, the first $\Delta$ arises from the difference between folded and unfolded states (e.g., $\Delta E=\langle E_{u}\rangle-\langle E_{f}\rangle$), and the second $\Delta$ arises from the free energy difference between arginine solution ($\Delta E_{arg}$) and water ($\Delta E_{wat}$) (e.g., $\Delta\Delta E=\Delta E_{arg}-\Delta E_{wat}$). With increasing arginine concentration, we observed an increasingly favorable cavitation component (Fig. 1h). Lin and Timasheff 79 previously connected protein stability with cavity formation and vapor-liquid surface tension. They proposed that the expansion of a protein-containing cavity requires more energy in the presence of an additive that increases surface tension at the cavity-solution interface. Experimentally, increasing arginine concentration has been observed to increase the vapor-liquid surface tension of aqueous solutions80, which may explain the dependence of $\Delta\Delta G_{cav}$ observed in the present study.
Looking at the trends in $\Delta\Delta E_{pw}$, we found that in 0.25 M arginine solutions, the polymer-water component favors polymer unfolding relative to in pure water (Fig. 1h). At 0.5 M and 1.0 M arginine concentrations, this component favors polymer collapse. With arginine present, the local domain of the polymer exhibits a reduction in the average number of water molecules (Fig. S10), indicating an effective expulsion of water. This, in turn, diminishes polymer-water interactions that resist polymer collapse.
The polymer-arginine contribution is observed to favor the folded polymer state at 0.25 M and 0.5 M arginine concentrations relative to in pure water (Fig. 1h). However, at 1.0 M arginine concentrations, polymer-arginine interactions promote polymer unfolding. The fact that high concentrations of arginine favored polymer unfolding is evidence that the direct mechanism of arginine-induced stabilization cannot fully explain the effect of this additive on hydrophobic polymer collapse.
To investigate potential competing effects of direct and indirect mechanisms, we combined the components of our PMF decomposition, delineating between those linked to direct effects (polymer-arginine and polymer-chloride; $\Delta\Delta G_{dir}$) and indirect effects (cavitation and polymer-water; $\Delta\Delta G_{ind}$) of arginine (Fig. 2a).
Figure 2: Contributions to the free energy of hydrophobic polymer unfolding in 0.25 M, 0.50 M, and 1.0 M (a) arginine, (b) guanidinium, and (c) glycine solutions. Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), direct interactions ($\Delta\Delta G_{dir}$), and indirect interactions ($\Delta\Delta G_{ind}$) and polymer are shown. Increasing additive concentration is denoted by increased shading (light to dark; left to right). Mean values are reported from three replicate REUS simulations. Error bars were estimated via error propagation (see SI for details).
We discovered that, with increasing concentration, the mechanism underlying the effects of arginine transitions from direct to indirect dominance. At 0.25 M, cavity formation and polymer-water interactions oppose polymer collapse, while arginine-polymer interactions favor collapse. The balance of these components gives rise to $\Delta\Delta G_{dir}>\Delta\Delta G_{ind}$, resulting in net stabilization of folded conformations and supporting the direct mechanism hypothesis (Fig. 2a).
In contrast, for the high-concentration regime (0.5 M and 1.0 M), cavity formation and polymer-water attractive interactions favor polymer collapse, while attractive arginine-polymer interactions favor extension of the hydrophobic polymer. In this case, indirect components dominated the free energy difference ($\Delta\Delta G_{dir}<\Delta\Delta G_{ind}$), stabilizing the polymer collapse and supporting the indirect hypothesis (Fig. 2a).
Thus, within the range of concentrations studied, we have uncovered that arginine exists at the edge of a mechanistic flip between direct- and indirect-dominated stabilization of many-body hydrophobic interactions. The identification of this mechanistic switch may explain the wide variety of hypotheses in the existing arginine literature. Because arginine is situated on this razor’s edge, small changes associated with the chemistry of a protein surface, the addition of cosolvents to solution, or differences in sample preparation, may cause significant changes in the modulation of hydrophobic interactions due to arginine.
Arginine is comprised of a polar backbone and an aliphatic side chain characterized by a guanidinium group. To investigate the roles of these components on hydrophobic polymer collapse, we completed an additional PMF decomposition in guanidinium and glycine solutions (see SI for simulation details). At all concentrations under study, we observed that guanidinium favors the hydrophobic polymer collapse primarily via a direct mechanism, while glycine stabilizes polymer collapse primarily via an indirect mechanism (Fig. 2b,c).
In the case of guanidinium, stabilization is driven entirely by attractive polymer-guanidinium interactions that favor collapse, while polymer-water interactions and cavity formation oppose polymer folding (Fig. 2b). In glycine solutions, however, stability is driven by the inverse mechanism; polymer-water interactions and the cavitation component favor collapse, while folding is opposed by attractive polymer-glycine interactions (Fig. 2c). Based on these findings, we characterize arginine as exhibiting a guanidinium-like mechanism at low concentrations and a glycine-like mechanism at high concentrations.
While glycine is known to be an effective stabilizer of proteins,81, 82, 18 our observations obtained for guanidinium are somewhat surprising due to its common role as a protein denaturant 83, 84, 85. Several studies have stressed the importance of direct interactions in guanidinium-induced denaturation, primarily via breaking salt bridges, competing for intra-protein hydrogen bonds, and interacting with aromatic moieties via cation-pi stacking.84, 86, 87. Usually, this occurs at high concentrations of guanidinium salts. Our findings suggest that while guanidinium may stabilize hydrophobic interactions at low concentrations, this is outweighed by the denaturation mechanisms at high concentrations.
Thermodynamic analyses of arginine effects on polymer collapse discussed above indicate that direct and indirect mechanisms co-exist and compete at all concentrations under study. To probe this further, we characterize the molecular interactions between arginine, water, and the polymer. Specifically, we look at hydrogen bonding between arginine and water to characterize arginine-water interactions. Preferential interactions are used to elucidate polymer-arginine-water interactions.
[…]
$C(\tau)=\left\langle\frac{\sum_{i,j}{h_{ij}\left(t_{0}\right)}{h_{ij}\left(t_{0}+\tau\right)}}{\sum_{i,j}{h_{ij}\left(t_{0}\right)^{2}}}\right\rangle$ (3)
where $h_{ij}\left(t_{0}\right)$ is equal to 1 if there is a hydrogen bond between groups $i$ and $j$ at time $t_{0}$, and 0 if no hydrogen bond is present. An average over all possible values of the time origin $\mathrm{t_{0}}$ was taken over the last 5 ns of 20 ns equilibrium simulations.
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (${}^{-}$, NH+${}_{3}^{+}$) and the guanidinium (${}^{+}$) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, ${}^{-}$-water, ${}^{+}$-water, and NH+${}_{3}^{+}$-water. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the ${}^{+}$ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and ${}^{-}$ are shaded in blue and purple, respectively.
Overall, the number of backbone-water hydrogen bonds was observed to be greater than sidechain-water hydrogen bonds (Fig. 3a). We further observed the fraction of occupied hydrogen bonding sites to be higher for backbone groups than the sidechain (Fig. S7). The hydrogen bond existence autocorrelation function for the 0.25 M arginine solution revealed hydrogen bonds formed between backbone-water atoms are, on average, longer-lived than sidechain-water hydrogen bonds (Fig. 3b). As arginine concentration was increased, hydrogen bond lifetimes were also observed to increase for both arginine-water and water-water interactions (Fig. S8).
Pairwise radial distribution functions (RDFs) were computed between the water oxygen (OW) and either the alpha carbon (CA) or guanidinium carbon (CZ) of arginine to quantify the local structure of water around arginine molecules (Fig. 3c). The first peak in the OW-CA RDF was observed to increase slightly with concentration. This indicates preferential hydration of the backbone as more arginine molecules are introduced to the solution. There is, however, no such change observed in the OW-CZ RDF with concentration. A representative snapshot of the hydration shell of a single arginine molecule is shown in Fig. 3d (2D representation is shown in Fig. S9. Together, these results indicate that the backbone of arginine interacts favorably with water.
While we found arginine preferentially interacts with water via its backbone, we hypothesized arginine interacts with the polymer via its sidechain. It has been reported elsewhere that dehydration of the guanidinium face is important in forming face-face stacking interactions in aqueous guanidinium solutions.92 In our case, the dehydrated face of guanidinium is expected to play a key role in direct arginine-polymer interactions, similar to interactions observed between guanidinium and hydrophobic/aromatic protein residues.93, 7, 20, 19
Distribution of arginine with respect to the polymer is described via the preferential interaction coefficient ($\Gamma_{PA}$),94, 95, 96
$\Gamma_{PA}=-\left(\frac{\partial\mu_{P}}{\partial\mu_{A}}\right)_{m_{P},T,P}=\left(\frac{\partial m_{A}}{\partial m_{P}}\right)_{\mu_{A},T,P}$ (4)
where $\mu$ is the chemical potential, $m$ is the concentration and $W$, $P$, and $A$ refer to water, polymer, and the additive, respectively. This parameter is calculated in simulations using the two-domain formula97, 98, 99 given by,
$\Gamma_{PA}=\left\langle N_{A}^{\text{local }}-\left(\frac{N_{A}^{\text{bulk }}}{N_{W}^{\text{bulk }}}\right)N_{W}^{\text{local }}\right\rangle$ (5)
where $N$ represents the number of molecules of a given species and angular brackets denote an ensemble average. The local and bulk domain was separated by a cutoff distance $R_{cut}$ from the polymer. $\Gamma_{PA}$ gives a measure of the relative accumulation or depletion of an additive in the local domain of the hydrophobic polymer, with $\Gamma_{PA}>0$ indicating relative accumulation (preferential interaction) and $\Gamma_{PA}<0$ indicating relative depletion (preferential exclusion).
Wyman-Tanford theory relates the dependence of any equilibrium process (such as protein folding) and preferential interaction as:100, 101, 102
$-\left(\frac{\partial\Delta G^{u}}{\partial\mu_{A}}\right)=\Gamma_{PA}^{u}-\Gamma_{PA}^{f}$ (6)
As a result, denaturants are expected to have a greater preferential interaction coefficient in the unfolded ensemble, while stabilizing osmolytes have a greater preferential interaction coefficient in the folded ensemble41, 58, 103, 104. In 0.25 M arginine concentration, we observed greater preferential interactions with the folded state relative to the unfolded state (Fig. 4a; Fig. S11a-c). With increasing concentration, we observe a diminishing difference between $\Gamma_{PA}^{u}$ and $\Gamma_{PA}^{f}$, in line with the mechanistic flip from a direct- to indirect-dominated stabilization mechanism.
Figure 4: Preferential interaction coefficient values as a function of the cut-off distance for the local domain of the hydrophobic polymer for (a) arginine, (b) guanidinium, and (c) glycine. Dashed lines indicate values for the unfolded state, while solid lines denote the folded state. Increasing concentration is denoted by increased shading (light to dark). Mean values and errors were estimated from three replicate simulations. Errors are reported as standard deviations from mean values.
Preferential interaction coefficients for guanidinium and glycine solutions are shown in Fig. 4b and c, respectively. We observed that at all concentrations, guanidinium preferentially interacts with the hydrophobic polymer (Fig. 4b). This is consistent with experimental evidence, as well as a prior simulation study that observed attractive guanidinium-polymer interactions for a model hydrophobic polymer39, 70. Glycine, meanwhile, was found to be preferentially excluded from the local domain of the hydrophobic polymer (Fig. 4c). This finding is consistent with the observed preference for the backbone of arginine to hydrogen bond with water, relative to the sidechain. Elsewhere, glycine has been observed to deplete from the surface of several model miniproteins, consistent with our findings 104.
To explore whether the preferential interactions of arginine with polymer and water are accompanied by preferential orientations, we computed an orientation parameter inspired by Shukla and Trout.105 This parameter was computed as the angle formed in three-dimensional space between polymer-CZ and CZ-CA vectors (Fig. 5a). The monomer closest to CZ is taken for the polymer-CZ vector. Angles where $\theta>90\degree$ indicate the arginine backbone orients towards the bulk solvent, while $\theta<90\degree$ indicates the arginine backbone orients towards the polymer. We observed that for all concentrations, the probability $P(\theta)$ is skewed towards angles greater than 90°$\degree$ (Fig. 5b-d).
Further, for all arginine concentrations, the mean value of $\theta$ for the folded state is greater than that observed for unfolded configurations (Fig. 5b-d). This preferential orientation of arginine enables the hydrophobic face of the guanidinium sidechain to interact with the hydrophobic polymer, while extension of the backbone towards the bulk enables additional interactions with either water or other free arginine molecules. The greater ability for arginine to adopt preferred orientations in the folded state, particularly at 0.25 M, may partially explain the favorable $\Gamma_{PA}$ values described previously.
Figure 5: Preferential orientation of arginine relative to the hydrophobic polymer. (a) Representation of the three-body angle, $\theta_{Pol-CZ-CA}$. $P(\theta)$ is shown for (b) 0.25 M, (c) 0.5 M, and (d) 1.0 M arginine concentrations. Solid lines denote the probability distribution for folded conformations, while dashed lines indicate the unfolded state. Vertical lines represent the ensemble average of each state.
Overall, our findings illuminate the intricate mechanisms underlying the multi-faceted effects of arginine on hydrophobic polymer collapse. Arginine was observed to increase the favorability of many-body hydrophobic interactions, a key factor in protein stabilization. Our observations reveal a nuanced interplay in the impact of arginine on hydrophobic interactions, teetering on the edge of a mechanistic flip. At low concentrations, direct sidechain-driven interactions dominate, shifting to indirect backbone-driven effects at high concentrations.
The simultaneous presence of competing direct and indirect effects implies that changes in the chemistry of a protein surface, the addition of co-additives to solution, or differences in sample preparation may cause significant changes in the mechanism of action of arginine. A shift towards the direct mechanism risks denaturing native proteins by disrupting electrostatic and hydrogen-bonding interactions. Conversely, a shift towards the indirect mechanism may yield stabilization of native proteins through preferential hydration. Our results suggest that arginine is uniquely situated for use in formulations due to its potential tunable, context-dependent properties. Hence, while arginine may not be considered the universal stabilizer it once was, its balance between direct- and indirect-driven stabilization of hydrophobic interactions solidifies its significance in formulation design.
{acknowledgement}
This material is based upon work supported by the National Science Foundation under DMREF Grant Nos. 2118788, 2118693, and 2118638.
{suppinfo}
Additional simulation details, error estimations, and figures are located in the Supporting Information.