### v3

# Flipping Out: Role of Arginine in Hydrophobic Interactions and Biological Formulation Design

Abstract
Arginine has been a mainstay in biological formulation development for decades. To date, the way arginine modulates protein stability has been widely studied and debated. Here, we employed a hydrophobic polymer to decouple hydrophobic effects from other interactions relevant to protein folding. While existing hypotheses for the effects of arginine can generally be categorized as either direct or indirect, our results indicate that direct and indirect mechanisms of arginine co-exist and oppose each other. At low concentrations, arginine was observed to stabilize hydrophobic polymer collapse via a sidechain-dominated direct mechanism, while at high concentrations, arginine stabilized polymer collapse via a backbone-dominated indirect mechanism. When adding partial charges to sites on the polymer, arginine destabilized polymer collapse. Further, we found arginine-induced destabilization of a model virus similar to direct-mechanism destabilization of the charged polymer, and concentration-dependent stabilization of a model protein similar to the indirect mechanism of hydrophobic polymer stabilization. These findings highlight the modular nature of the widely used additive arginine, with relevance in the design of stable biological formulations.
MTU] Department of Chemical Engineering, Michigan Technological University, Houghton, MI 49931, USA UMA] Department of Chemical Engineering, University of Massachusetts Amherst, MA 01003, USA MTU] Department of Chemical Engineering, Michigan Technological University, Houghton, MI 49931, USA UMA] Department of Chemical Engineering, University of Massachusetts Amherst, MA 01003, USA
1 Introduction
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification.7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation,10, 11, 12 denaturation,13, 14, 15 and inactivation16, 17 of certain proteins and viruses. Additional studies have found that arginine mechanisms are dependent on concentration.18, 19, 20, 21 Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
In general, additive effects on protein stability are thought to be either direct or indirect. Direct mechanisms involve direct protein-additive interactions, while indirect mechanisms influence protein stability by modulating the surrounding solvent structure. It remains debated whether arginine acts primarily via a direct or indirect mechanism. Several studies called attention to direct interactions between arginine and aromatic residues,7, 22, 23 acidic residues,12, 20, 24 and hydrophobic moieties.25, 26 Other studies have proposed clusters of free arginine molecules in solution enable the crowding out of protein-protein interactions,27, 23, 28, 20 or alteration of hydration shell water dynamics.29, 30 The wide range of observations related to the role of arginine on protein stability suggests that arginine harbors diverse, context-dependent mechanisms.
To elucidate the mechanisms through which arginine influences protein stability, this study focuses on its effects on hydrophobic interactions. Hydrophobic interactions are key in several biologically relevant phenomena, including protein folding and stability.31, 32, 33, 34, 35, 36, 37, 38, 39 Additives in solutions are known to modulate the strength of hydrophobic interactions, and in turn, the stability of proteins.40, 41, 42, 43, 44, 45, 46 For example, simulation studies have shown trimethylamine N-oxide (TMAO) has a negligible effect on or strengthens hydrophobic interactions,42, 47, 48, 49, 50, 51 while these interactions are weakened in urea solutions.52, 53, 41, 54, 55, 43, 44, 56 Indeed, the effects of TMAO and urea on hydrophobic interactions are consistent with their experimentally-observed roles as a protein stabilizer and denaturant, respectively.57, 58, 59, 45
Molecular dynamics (MD) simulations have provided valuable insights towards understanding these mechanisms as they relate to hydrophobicity. Several studies have highlighted the utilization of a hydrophobic polymer model for describing the role of solvent and additives on protein-like collapse. 60, 44, 61, 62, 43, 63 The use of a hydrophobic polymer model enables the decoupling of additive effects on hydrophobic vs other interactions, which is challenging in experiments. Additionally, comparison of arginine mechanisms on purely hydrophobic interactions allows indirect insights into experimentally-observed effects. For example, studies in which effects mirror what we observe from the hydrophobic polymer are likely to reflect a dominant mediation of hydrophobic interactions. In the present study, we utilize MD simulations of a hydrophobic polymer to characterize the effects of arginine on many-body hydrophobic interactions pertinent to protein stability, and contextualize these findings within larger-scale models for biological formulations.
Overall, we found arginine stabilizes hydrophobic polymer collapse at all concentrations under study. Interestingly, we discovered arginine sits on the edge of a mechanistic flip, balancing between direct- and indirect-dominated effects. As a consequence of this balance, we found subtle modulation of the polymer chemistry (via partial charge incorporation) changes arginine from a stabilizing additive to a destabilizer of polymer collapse. Consequently, in practical examples of formulation design, we observed arginine has variable effects on a model virus and protein.
2 Methods
2.1 Hydrophobic Polymer System Setup and Molecular Dynamics Simulations
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). All simulations were performed using GROMACS 2021.464, 65 with the PLUMED 2.8.0 66, 67 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$42 Box dimensions were defined such that 1.5 nm of space separated the fully elongated polymer from the nearest box edge. For simulations with no polymer, the same box dimensions were used. Arginine was modeled in accordance with a pH of 7, resulting in a protonated sidechain. An equal number of arginine molecules and Cl- atoms were added to the box until the desired concentration was reached. The TIP4P/200568 model was used to describe water, and the CHARM22 force field was used for arginine and Cl-.69
All simulations were initially subject to energy minimization using the steepest descent algorithm. NVT equilibration was carried out for 1 ns at 300 K, followed by a 1 ns NPT equilibration at 300 K and 1 atm. During equilibration, temperature was controlled according to the V-rescale thermostat,70 while pressure was controlled via the Berendsen barostat.71 Following equilibration, NPT production runs were completed using the Nosé-Hoover thermostat72 and Parrinello-Rahman barostat.73 Production runs were completed for 20 ns for arginine/water systems, and between 50-250 ns per window for arginine/polymer/water replica exchange umbrella sampling (REUS) runs (Table 2). In all simulations, the Particle Mesh Ewald (PME) algorithm was used for electrostatic interactions with a cut-off of 1 $\mathrm{nm}$. A reciprocal grid of 42 x 42 x 42 cells was used with $\mathrm{4^{th}}$ order B-spline interpolation. A single cut-off of 1 $\mathrm{nm}$ was used for van der Waals interactions. The neighbor search was performed every 10 steps. Lorentz-Berthelot mixing rules74, 75 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
2.2 Replica Exchange Umbrella Sampling
REUS76 simulations were completed to sample the hydrophobic polymer conformational landscape in arginine solutions. The radius of gyration ($R_{g}$) of the hydrophobic polymer was used as the reaction coordinate to describe polymer folding/unfolding. 12 umbrella potential windows were centered between $R_{g}$ = 0.3 and $R_{g}$ = 0.9 nm, with a spacing of 0.05 nm. A force constant of $K$ = 5000 kJ/mol/nm2 was used in all windows, with the exception of the window centered at $R_{g}$ = 0.45, which used $K$ = 1000 kJ/mol/nm2 (see SI for details).
The potential of mean force (PMF) along the radius of gyration of the polymer was calculated as ${W(R_{g})=-k_{B}Tln(P(R_{g}))}$. Biased probability distributions were reweighted according to the Weighted Histogram Analysis Method (WHAM).77 The free energy of polymer unfolding ($\mathrm{\Delta G_{u}}$) was calculated according to:
[…]
where $\mathrm{R_{g,cut}}$ was determined as the point between the folded and unfolded states where $\frac{\partial W(R_{g})}{\partial R_{g}}=0$.
We decomposed the PMF into individual components to further investigate the role of arginine in polymer collapse. Following the methods outlined by several others,78, 79, 46, 80 the PMF was decomposed as:
$W(R_{g})=W_{vac}(R_{g})+W_{cav}(R_{g})+E_{pw}(R_{g})+E_{pa}(R_{g})+E_{pc}(R_{g})$ (2)
$W_{vac}(R_{g})$ captures intrapolymer degrees of freedom and was obtained from independent REUS simulations of the polymer in vacuum. $E_{pw}(R_{g})$, $E_{pa}(R_{g})$, and $E_{pc}(R_{g})$ are average polymer-water, polymer-arginine, and polymer-chloride interaction energies, respectively. The remaining term is $W_{cav}(R_{g})$, which is the cavitation component and quantifies the energetic cost of forming a cavity – of the same size and shape as the polymer – in the solution.
2.3 Preferential Interaction Coefficients
Distribution of arginine with respect to the polymer can be described via the preferential interaction coefficient ($\Gamma_{PA}$),81, 82, 83
$\Gamma_{PA}=-\left(\frac{\partial\mu_{P}}{\partial\mu_{A}}\right)_{m_{P},T,P}=\left(\frac{\partial m_{A}}{\partial m_{P}}\right)_{\mu_{A},T,P}$ (3)
where $\mu$ is the chemical potential, $m$ is the concentration and $W$, $P$, and $A$ refer to water, polymer, and an additive, respectively. This parameter is calculated in simulations using the two-domain formula84, 85, 86 given by:
$\Gamma_{PA}=\left\langle N_{A}^{\text{local }}-\left(\frac{N_{A}^{\text{bulk }}}{N_{W}^{\text{bulk }}}\right)N_{W}^{\text{local }}\right\rangle$ (4)
where $N$ represents the number of molecules of a given species and angular brackets denote an ensemble average. The local and bulk domain was separated by a cutoff distance $R_{cut}$ from the polymer. $\Gamma_{PA}$ gives a measure of the relative accumulation or depletion of an additive in the local domain of the hydrophobic polymer, with $\Gamma_{PA}>0$ indicating relative accumulation (preferential interaction) and $\Gamma_{PA}<0$ indicating relative depletion (preferential exclusion).
2.4 Hydrogen Bond Analysis
Hydrogen bonds were calculated according to geometric criteria of a donor-acceptor distance of $r\leq$ 0.35 $\mathrm{nm}$ and donor-hydrogen-acceptor angle within 150° < $\theta$ < 180°.87 Hydrogen bond existence correlation functions for water-water and arginine-water interactions were estimated according to:88, 89, 90
$C(\tau)=\left\langle\frac{\sum_{i,j}{h_{ij}\left(t_{0}\right)}{h_{ij}\left(t_{0}+\tau\right)}}{\sum_{i,j}{h_{ij}\left(t_{0}\right)^{2}}}\right\rangle$ (5)
where $h_{ij}\left(t_{0}\right)$ is equal to 1 if there is a hydrogen bond between groups $i$ and $j$ at time $t_{0}$, and 0 if no hydrogen bond is present. An average over all possible values of the time origin $t_{0}$ was taken over the last 5 ns production simulations.
2.5 Contact Coefficients
Contact coefficients ($CC$) give a measure of excipient preference for interacting with specific residues. Here, we computed $CC$s as described by Stumpe and Grubmüller91:
$CC_{X}=\frac{N_{X-A}}{N_{X-W}}\frac{M_{W}}{M_{A}}$ (6)
where $N_{X-A}$ and $N_{X-W}$ are the number of atomic contacts of protein residue $X$ with an additive or water molecules, respectively. Atoms were defined to be in contact if any pair of heavy atoms were within a 0.35 nm cut-off. $CC$s are normalized by the total number of additive atoms ($M_{A}$) and water atoms ($M_{W}$) in solution. Values of $CC_{X}~{}\textgreater~{}1$ indicate preferential interaction with an additive for residue $X$, while $CC_{X}~{}\textless~{}1$ denotes preferential interaction with water.
2.6 Protein and Virus Simulation Setup
Porcine parvovirus (PPV) and hen egg white lysozyme (HEWL) were used as large-scale models for formulation design. To model PPV, we constructed a surface model using 15 monomers from the published crystal structure (PDB: 1K3V).92 This was achieved by selecting three reference $C_{\alpha}$ atoms from the 5-fold pore of the viral surface, and aligning vector normal to these points with the z-axis of the simulation box. Box dimensions were 24 x 24 x 12 nm, with the capsid surface model extending approximately 9 nm in the z-direction. Arginine molecules were added to the exterior of the capsid surface to reach a target concentration of 0.25 M, and NaCl atoms were distributed throughout the box to both neutralize the system and reach a target concentration of 0.15 mM. To prevent diffusion of arginine molecules into the capsid interior, a wall was placed at z = 0 nm, and the capsid surface was positioned on top of this wall (Fig. S12). A position restraint with a force constant of 1000 kJ/mol/nm2 was applied to all capsid atoms within 1.5 nm from the wall. Capsid atoms between 1.5 nm and 3.0 nm from the wall were restrained with a force constant between 1-1000 kJ/mol/nm2, with atoms further from the wall scaled to a lesser extent. Capsid atoms greater than 3.0 nm from the wall were left fully flexible. PPV simulations were carried out in the NVT ensemble – first for 1 ns to reach the target temperature, followed 50 ns to allow the solvent to relax. 100 ns production runs were used for further analysis.
HEWL was modeled from an available crystal structure (PDB: 1LYZ),93 and simulated in solution with 0.1, 0.2, 1.0, and 2.0 M arginine. HEWL simulations were carried out in the NVT ensemble for 1 ns to equilibrate to the target temperature, followed by 1 ns NPT simulations to reach the target density. HEWL production runs were completed for 100 ns in the NPT ensemble. For both HEWL and PPV, the force field, thermostats, and barostats matched those used for the hydrophobic polymer systems.
2.7 Experimental Temperature Stability Studies
Liquid samples of PPV in arginine solutions were prepared in triplicates and were put either in a heat block at 60 °C 94 or in a fridge at 4 °C as the control samples. 72 hours later, the titer of PPV was determined by 3-(4,5-dimethylthiazol-2-yl)-2,5-diphenyltetrazolium bromide (MTT) colorimetric cell viability assay.95 Differential scanning fluorimetry (DSF) was performed to compute the hydrophobic exposure temperature (HET) of HEWL in arginine solutions.96, 97 More details for both experimental methods are included in the supporting information.
3 Results and Discussion
3.1 Arginine favors hydrophobic polymer collapse
Fig. 1a shows the PMF along the $R_{g}$ reaction coordinate. In all solutions, free energy minima were observed at approximately 0.4 and 0.8 nm (configurations labeled I and III in Fig. 1g), along with a prominent free energy barrier at $\sim$0.6 nm representing the transition between folded and unfolded states (Fig. 1a). In pure water, hydrophobic collapse is unfavorable, with the unfolded state favored by $\sim$0.3 kT. In contrast, at all arginine concentrations, the folded state of the polymer is favored relative to pure water, and a monotonic increase in $\Delta G_{u}$ is observed (Fig. S4). An additional barrier at $\sim$0.45 nm was identified separating two folded states, labeled as I and II in Fig. 1a.
The folded polymer ensemble in arginine solutions exhibits free energy minima corresponding to globular ($\sim$0.4 nm) and hairpin-like ($\sim$0.5 nm) configurations (labeled I and II in Fig. 1g). Arginine clusters encapsulating the hydrophobic polymer are observed in each state (Fig. S6). We propose that the free energy barrier separating these two states arises from an energetic penalty associated with breaking these encapsulating clusters. Such a mechanism is similar to that observed by Li et al.,26 who observed arginine-mediated suppression of hydrophobic association.
Figure 1: PMF decomposition in 0.0 M, 0.25 M, 0.5 M, and 1.0 M arginine solutions. (a) The PMF obtained along the $R_{g}$ reaction coordinate, $W(R_{g})$, (b) vacuum component, $W_{vac}$, (c) cavitation component, $W_{cav}$, (d) polymer-water interactions, $E_{pw}$, (e) polymer-arginine interactions, $E_{pa}$, and (f) polymer-chloride interactions, $E_{pc}$. (g) Representative configurations along the reaction coordinate as denoted in (a) as I, II, and III. (h) Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), cavitation contribution ($\Delta\Delta G_{cav}$), polymer-water interactions ($\Delta\Delta E_{pw}$), and polymer-arginine interactions ($\Delta\Delta E_{pa}$). Increasing arginine concentration is denoted by increased shading (light to dark) and is indicated by arrows in (a-f). The polymer in water alone is shown in black, where appropriate. Mean values were estimated from three replicate REUS simulations. Error bars are reported as described in the SI. All plots are normalized to 0 at $R_{g}=0.4$ nm, where appropriate.
3.2 Decomposition of arginine effects on polymer folding
Fig. 1b-f shows the decomposition of the PMF into various components. The vacuum component, $W_{vac}$, favors the folded state of the polymer, associated with favorable intrapolymer interactions and configurational entropy upon collapse (Fig. 1b). Because this component does not depend on the presence of arginine, a balance of the remaining components dictates the effect of arginine on hydrophobic polymer collapse.
Near large idealized solutes or hydrophobic interfaces, water dewets and forms a vapor-liquid-like interface.98, 99 Similar behavior may be expected for our hydrophobic polymer; hence, we computed $W_{cav}$, which depends on both the size and shape of the polymer and is related to the vapor-liquid surface tension.79, 100, 78 The cavitation component favors the folded state (Fig. 1c), reflecting a strong hydrophobic driving force for polymer collapse.60
Attractive polymer-water interactions become more favorable with increasing $R_{g}$ (Fig. 1d), indicating polymer-water interactions oppose polymer collapse. It is worth noting that the free energy minima at $\sim$0.8 nm in the unfolded ensemble is observed in an aqueous environment (Fig. 1a), but not in vacuum (Fig. 1b). This minima arises due to favorable polymer-water interactions, consistent with prior MD simulations that showed water-mediated interactions drive large hydrophobic solutes apart.101, 102, 61 Additionally, sufficient dewetting of the hydrophobic polymer is a known bottleneck to collapse,103, 60, 78 resulting in the free energy barrier at $\sim$0.6 nm separating the folded and unfolded states.
Attractive polymer-arginine interactions approach an energetic minima at $\sim$0.5 nm (Fig. 1e), giving rise to the global minimum observed in the overall PMF. Polymer-Cl- interactions were observed on the order of thermal fluctuations, consistent with previous observations that Cl- ions are depleted from the local domain of hydrophobic solutes.104, 41
Fig. 1h shows the change in each component upon unfolding in arginine solution relative to that observed in water. In Fig. 1h, the first $\Delta$ arises from the difference between folded and unfolded states (e.g., $\Delta E=\langle E_{u}\rangle-\langle E_{f}\rangle$), while the second $\Delta$ arises from the free energy difference between arginine solution ($\Delta E_{arg}$) and water ($\Delta E_{wat}$) (e.g., $\Delta\Delta E=\Delta E_{arg}-\Delta E_{wat}$). With increasing arginine concentration, we observed an increasingly favorable cavitation component for folding (Fig. 1h). Lin and Timasheff 105 previously connected protein stability with cavity formation and vapor-liquid surface tension. They proposed that the expansion of a protein-containing cavity requires more energy in the presence of an additive that increases surface tension at the cavity-solution interface. Experimentally, increasing arginine concentration has been observed to increase the vapor-liquid surface tension of aqueous solutions,106 which may explain the dependence of $\Delta\Delta G_{cav}$ observed in the present study.
Looking at the trends in $\Delta\Delta E_{pw}$, we found that in 0.25 M arginine solutions, the polymer-water component favors polymer unfolding relative to in pure water (Fig. 1h). At 0.5 M and 1.0 M arginine concentrations, this component favors polymer collapse. With arginine present, the local domain of the polymer exhibits a reduction in the average number of water molecules (Fig. S10), indicating an effective expulsion of water. This, in turn, diminishes polymer-water interactions that resist polymer collapse.
The polymer-arginine contribution favors the folded polymer state at 0.25 M and 0.5 M arginine concentrations relative to in pure water (Fig. 1h). However, at 1.0 M arginine concentrations, polymer-arginine interactions promote polymer unfolding. Together, these results indicate that neither direct nor indirect mechanistic hypotheses alone can describe the effects of arginine on hydrophobic polymer folding.
3.3 Mechanistic flip with increasing arginine concentration
To investigate potential competing effects of direct and indirect mechanisms, we combined the components of our PMF decomposition, delineating between those linked to direct effects (polymer-arginine and polymer-Cl-; $\Delta\Delta G_{dir}$) and indirect effects (cavitation and polymer-water; $\Delta\Delta G_{ind}$) of arginine (Fig. 2a).
Figure 2: Contributions to the free energy of hydrophobic polymer unfolding in 0.25 M, 0.50 M, and 1.0 M (a) arginine, (b) guanidinium, and (c) glycine solutions. Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), direct interactions ($\Delta\Delta G_{dir}$), and indirect interactions ($\Delta\Delta G_{ind}$) and polymer are shown. Increasing additive concentration is denoted by increased shading (light to dark; left to right). Mean values are reported from three replicate REUS simulations. Error bars were estimated via error propagation (see SI for details).
We discovered that, with increasing concentration, the mechanism underlying the effects of arginine transitions from direct to indirect dominance. At 0.25 M, cavity formation and polymer-water interactions oppose polymer collapse, while arginine-polymer interactions favor collapse. The balance of these components gives rise to $\Delta\Delta G_{dir}>\Delta\Delta G_{ind}$, resulting in the net stabilization of folded conformations and supporting the direct mechanism hypothesis (Fig. 2a).
In contrast, for the high-concentration regime (0.5 M and 1.0 M), cavity formation and polymer-water attractive interactions favor polymer collapse, while attractive arginine-polymer interactions favor extension of the hydrophobic polymer. In this case, indirect components dominate the free energy difference ($\Delta\Delta G_{dir}<\Delta\Delta G_{ind}$), stabilizing polymer collapse and supporting the indirect hypothesis (Fig. 2a).
Thus, within the range of concentrations studied, we have uncovered that arginine exists at the edge of a mechanistic flip between direct- and indirect-dominated stabilization of many-body hydrophobic interactions. The identification of this mechanistic switch may explain the wide variety of hypotheses in the existing arginine literature. Because arginine is situated on this razor’s edge, small changes associated with the chemistry of a protein surface, the addition of cosolvents to solution, or differences in sample preparation, may cause significant changes in the modulation of hydrophobic interactions due to arginine.
3.4 Distinct roles of arginine’s sidechain and backbone
Arginine is comprised of a polar backbone and an aliphatic sidechain characterized by a guanidinium group. To investigate the roles of these components on hydrophobic polymer collapse, we completed an additional PMF decomposition in guanidinium and glycine solutions (see SI for simulation details). At all concentrations under study, we observed that guanidinium favors the hydrophobic polymer collapse primarily via a direct mechanism, while glycine stabilizes polymer collapse primarily via an indirect mechanism (Fig. 2b,c).
In the case of guanidinium, stabilization is driven entirely by attractive polymer-guanidinium interactions that favor collapse, while polymer-water interactions and cavity formation oppose polymer folding (Fig. 2b). In glycine solutions, however, stability is driven by the inverse mechanism; polymer-water interactions and the cavitation component favor collapse, while folding is opposed by attractive polymer-glycine interactions (Fig. 2c). Based on these findings, we characterize arginine as exhibiting a guanidinium-like mechanism at low concentrations and a glycine-like mechanism at high concentrations.
While glycine is known to be an effective stabilizer of proteins,107, 108, 21 our observations obtained for guanidinium are somewhat surprising due to its common role as a protein denaturant.109, 110, 111 Several studies have stressed the importance of direct interactions in guanidinium-induced denaturation, primarily via breaking salt bridges, competing for intra-protein hydrogen bonds, and interacting with aromatic moieties via cation-pi stacking.110, 112, 113 Usually, this occurs at high concentrations of guanidinium salts. Our findings suggest that while guanidinium may stabilize hydrophobic interactions at low concentrations, this is outweighed by denaturing mechanisms at high concentrations.
3.5 Resolving polymer-arginine-water interactions
Thermodynamic analyses of arginine effects on polymer collapse discussed above indicate that direct and indirect mechanisms co-exist and compete at all concentrations under study. To probe this further, we characterize the molecular interactions between arginine, water, and the polymer. Specifically, we look at hydrogen bonding between arginine and water to characterize arginine-water interactions. Preferential interactions are used to elucidate the balance of polymer-arginine-water interactions.
To describe how arginine interacts with water, we considered hydrogen bonding interactions between water and backbone (COO-, NH+${}_{3}^{+}$) or sidechain (Gdm+) atoms of arginine. Overall, the number of backbone-water hydrogen bonds was observed to be greater than sidechain-water hydrogen bonds (Fig. 3a). We further observed the fraction of occupied hydrogen bonding sites to be higher for backbone groups than the sidechain (Fig. S7). The hydrogen bond existence autocorrelation function for the 0.25 M arginine solution revealed hydrogen bonds formed between backbone-water atoms are, on average, longer-lived than sidechain-water hydrogen bonds (Fig. 3b). As arginine concentration was increased, hydrogen bond lifetimes were also observed to increase for both arginine-water and water-water interactions (Fig. S8). In principle, this may be considered a stabilizing property of arginine, as a growing body of literature has supported the role of stabilizing osmolytes in increasing hydrogen bond lifetimes and reducing water dynamics.114, 115, 116, 117, 118, 119, 120
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (COO-, NH+${}_{3}^{+}$) and the guanidinium (Gdm+) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, COO--water, Gdm+-water, and NH+${}_{3}^{+}$-water in 0.25 M arginine solution. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the Gdm+ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and COO- are shaded in blue and purple, respectively.
Pairwise radial distribution functions (RDFs) were computed between the water oxygen (OW) and either the alpha carbon (CA) or guanidinium carbon (CZ) of arginine to quantify the local structure of water around arginine molecules (Fig. 3c). The first peak in the OW-CA RDF was observed to increase slightly with concentration. This indicates preferential hydration of the backbone group as more arginine molecules are introduced to the solution. There is, however, no such change observed in the OW-CZ RDF with concentration. A representative snapshot of water interactions with a single arginine molecule is shown in Fig. 3d (2D representation is shown in Fig. S9). Together, these results indicate that the backbone of arginine is the primary site for interaction with water.
While we found arginine preferentially interacts with water via its backbone, we hypothesized arginine interacts with the polymer via its sidechain. It has been reported elsewhere that dehydration of the planar guanidinium face is important in forming face-face stacking interactions in aqueous guanidinium solutions.121 In our case, the dehydrated face of guanidinium is expected to play a key role in direct arginine-polymer interactions, similar to interactions observed between guanidinium and hydrophobic/aromatic protein residues.122, 7, 23, 22
Wyman-Tanford theory relates the dependence of any equilibrium process (such as protein folding) and preferential interaction as:123, 124, 125
$-\left(\frac{\partial\Delta G_{u}}{\partial\mu_{A}}\right)=\Gamma_{PA}^{u}-\Gamma_{PA}^{f}$ (7)
where $\Gamma_{PA}^{u}$ represents the preferential interaction coefficient in the unfolded state, while $\Gamma_{PA}^{f}$ represents the folded state. As a result, denaturants are expected to have a greater preferential interaction coefficient in the unfolded ensemble, while stabilizing osmolytes have a greater preferential interaction coefficient in the folded ensemble.45, 62, 126, 127 In 0.25 M arginine concentration, we observed greater preferential interactions with the folded state relative to the unfolded state (Fig. 4a; Fig. S11a-c). With increasing concentration, we observe a diminishing difference between $\Gamma_{PA}^{u}$ and $\Gamma_{PA}^{f}$, in line with the mechanistic flip from a direct- to indirect-dominated stabilization mechanism.
Figure 4: Preferential interaction coefficient values as a function of the cut-off distance for the local domain of the hydrophobic polymer for (a) arginine, (b) guanidinium, and (c) glycine. Dashed lines indicate values for the unfolded state, while solid lines indicate the folded state. Increasing concentration is denoted by increased shading (light to dark). Mean values and errors were estimated from three replicate simulations. Errors are reported as standard deviations from mean values.
Preferential interaction coefficients for guanidinium and glycine solutions are shown in Fig. 4b and c, respectively. We observed that at all concentrations, guanidinium preferentially interacts with the hydrophobic polymer (Fig. 4b). This is consistent with experimental evidence, as well as a prior simulation study that observed attractive guanidinium-polymer interactions with a model hydrophobic polymer.43, 79 Glycine, meanwhile, was found to be preferentially excluded from the local domain of the hydrophobic polymer (Fig. 4c). This finding is consistent with the observed preference for the backbone of arginine to hydrogen bond with water, relative to the sidechain. Elsewhere, glycine has been observed to deplete from the surface of several model miniproteins, consistent with our findings. 127
To explore whether the preferential interactions of arginine with polymer and water are accompanied by preferential orientations, we computed an orientation parameter inspired by Shukla and Trout.128 This parameter was computed as the angle formed in three-dimensional space between polymer-CZ and CZ-CA vectors (Fig. 5a). The monomer closest to CZ is taken for the polymer-CZ vector. Angles where $\theta>90\degree$ indicate the arginine backbone orients towards the bulk solvent, while $\theta<90\degree$ indicates the arginine backbone orients towards the polymer. We observed that for all concentrations, the probability $P(\theta)$ is skewed towards angles greater than 90°$\degree$ (Fig. 5b).
Further, for all arginine concentrations, the mean value of $\theta$ for the folded state is greater than that observed for unfolded configurations (Fig. 5b). This preferential orientation of arginine enables the hydrophobic face of the guanidinium sidechain to interact with the hydrophobic polymer, while extension of the backbone towards the bulk enables additional interactions with either water or other free arginine molecules. The greater ability for arginine to adopt preferred orientations in the folded state, particularly at 0.25 M, may partially explain the favorable $\Gamma_{PA}$ values described previously.
Figure 5: Preferential orientation of arginine relative to the hydrophobic polymer. (a) Representation of the three-body angle, $\theta$. (b) $P(\theta)$ is shown for 0.25 M, 0.5 M, and 1.0 M arginine concentrations. Solid lines denote the probability distribution for folded conformations, while dashed lines indicate the unfolded state. Increased concentration is denoted by increased shading.
4 Contextualizing the Mechanistic Flip of Arginine
4.1 Arginine destabilizes a charged polymer
To explore whether we could shift arginine further towards a direct-dominated mechanism, we modified our hydrophobic polymer to include four beads with opposing partial charges (Fig.  6a). In pure water, charged polymer collapse was observed to be favorable, resulting in a prominent free energy minimum at $R_{g}\sim$ 0.5 nm (Fig.  6b). Relative to the hydrophobic polymer, this is indicative of a conformational preference of the polymer to adopt hairpin-like, rather than globule-like, configurations (I and II in Fig.  6a, respectively).
Upon addition of either 0.25 M or 1.0 M arginine, folding of the charged polymer becomes less favorable (Fig.  6c). We attribute this to a preference for arginine to interact with the charged sites, which are more accessible to arginine in the unfolded state. Indeed, at both concentrations, attractive polymer-arginine interactions dominate (Fig.  6d), driving polymer unfolding. This model demonstrates that even subtle changes to the chemistry of a macromolecule can re-balance arginine mechanisms.
Figure 6: Charged polymer simulations. (a) Locations of the negative (blue) and positive (red) partial charges. (b) PMFs associated with the charged polymers. (c) Free energy of polymer collapse as a function of arginine concentration. (d) Change in PMF components in Arg solution relative to pure water, for the charged polymer. Increasing concentration is denoted by increased shading.
4.2 Implications for formulation design
Arginine-induced destabilization of a charged polymer illustrates that, when hydrophobic interactions compete with other effects, the effectiveness of arginine as a stabilizer can be altered. To better understand this feature of arginine, we explored the temperature stability of two models for formulation design: PPV and HEWL.
PPV is a non-enveloped virus with a single stranded DNA genome.129 The PPV capsid is a spherical shell comprised of 60 copies of viral proteins (VP) VP1, VP2, and VP3 in a 1:10:1 ratio, arranged in an icosahedral symmetry.92 Due to the relatively small size and structural simplicity of PPV, recent studies have employed the virus as a model for investigating virus purification and thermostabilization techniques.94, 130, 131, 95
HEWL, meanwhile, is a relatively small protein with a well-defined fold (PDB: 1LYZ) and high stability.93, 132 Due to these features, it has been widely used as a model for protein folding133, 134, 135 and exploring osmolyte effects.136, 137, 107 HEWL has also been used to investigate aggregation-suppressing effects of arginine. Several studies have proposed that arginine interacts favorably with aromatic and acidic residues of HEWL, which limits solvent exposure of aggregation-prone patches.7, 22, 23, 20, 138 Arginine has also been observed to enhance the heat-induced aggregation of bovine serum albumin and $\beta$-lactoglobulin, but not HEWL – highlighting its context-dependent effects.11
To understand the effect of arginine on these biomolecules, temperature stability assays were carried out at different concentrations of arginine. For PPV, we completed an infectivity assay (see SI for experimental details) following virus incubation at 60 °C for 3 days – sufficiently long to observe a significant decrease in the infectious titer.94 Log reduction values (LRV) describe the decrease in the infectious titer of heat-treated virus, relative to the initial virus solution. We found, at all investigated arginine concentrations, more infectious PPV is lost relative to in buffer alone, resulting in negative $\Delta$LRV values (Fig.  7a). Such a finding indicates reduced temperature stability of the virus in arginine solutions.
Figure 7: Temperature stability of (a) PPV and (b) HEWL as a function of arginine concentration. $\Delta$LRV is reported as LRVBuffer - LRVArg, while $\Delta$THE $=$ THE,Arg - THE,buffer. Increasing concentration is denoted by increased shading. Mean and standard deviation are estimated from triplicate measurements.
For HEWL, we completed a thermal shift assay (see SI for details) to quantify the hydrophobic exposure temperature ($T_{HE}$) of the protein.96, 97 At low concentrations (0.2-1.0 M) of arginine, a decrease in $T_{HE}$ is observed relative to buffer alone, indicating destabilization of HEWL (Fig.  7b). With increasing arginine concentration (1.5 M), the temperature stability of HEWL improves. Such a finding closely resembles the temperature stability of ovalbumin and lysozyme in arginine solutions reported by Vagenende et al.20 Additionally, 1.0 M arginine was found to increase protein interactions with hydrophobic chromatographic materials.20 This finding highlights that, when experimentally-observed effects of arginine mirror what we observed with the hydrophobic polymer, hydrophobic interactions are likely to dominate. On the other hand, our results contrast findings presented by Platts and Falconer,21 who observed temperature-stabilizing effects at low arginine concentrations and destabilization of HEWL at high concentrations.
4.3 Molecular-level investigation of formulation models
The lack of consensus regarding the effects of arginine – further accentuated by our investigation of PPV and HEWL temperature stability – may be rationalized by the positioning of arginine at the edge of the mechanistic flip described in the present work. Additionally, we have shown that when charged beads are added to a hydrophobic polymer, this subtle change results in complete arginine-induced destabilization. To better understand the connection between the mechanistic flip of arginine and the relevance of differing molecular contexts, we completed straightforward MD simulations of HEWL and PPV systems in arginine solution.
The PPV major capsid protein, VP2, is known to self-assemble into virus-like particles that are non-pathogenic, but are morphologically similar to native PPV virions.139 Experimental identification of this structure92 enables PPV for molecular-level investigations via molecular simulations. On the capsid surface, prominent structural features include a canyon surrounding a pore at the 5-fold axis of symmetry, protruding spikes located at the 3-fold axis of symmetry, and a dimple on the 2-fold axis of symmetry.92 We found that the use of 15 VP2 proteins is the minimum system size required to capture these structural features, enabling simulations at timescales accessible to MD without sacrificing atomistic resolution. Despite this truncation of the viral capsid, our resulting system size was $\sim$750,000 atoms.
At the PPV surface in 0.25 M arginine solution, we observed preferential arginine accumulation at various sites across the surface (Fig.  8a,c). Specifically, via calculation of contact coefficients of different residue types distributed across the PPV capsid surface, we identified significant accumulation of arginine near negatively charged glutamate and aspartate residues. From these findings, we hypothesize that direct interactions between arginine and charged residues at the PPV surface drive the instability observed from experiments. In this case, the dominating mechanism may be similar to that observed for arginine-induced unfolding of a charged polymer.
For HEWL, we completed simulations of 0.1 M, 0.2 M, 1.0 M, and 2.0 M arginine solutions. With increasing concentration, we observed an increase in preferential exclusion of arginine and reduction of arginine density local to the protein (Fig.  8b,d). This finding, along with the increase in HEWL melting temperature observed at high arginine concentration, suggests that arginine imparts stabilization of HEWL via indirect effects. On the other hand, direct arginine-HEWL interactions may be destabilizing, which explains the decreased melting temperature of HEWL while arginine is less excluded from the local domain of the protein.
Figure 8: Arginine distributions from PPV and HEWL systems. (a) Arginine contact coefficients at the PPV 5-fold surface. Bars are shown for different residue types, hydrophobic (green), hydrophilic (yellow), positive (blue), and negative (red), and was computed separately for backbone atoms (gray). (b) Preferential interaction coefficient for arginine/HEWL. Increasing concentration is denoted by increased shading (light to dark). (c) Representative snapshot and volumetric density plots of the (c) PPV 5-fold surface and (d) HEWL in arginine solutions. In c and d, PPV and HEWL are shown in a surface representation. Protein residues are colored according to residue type as in panel a. Arginine density at an isovalue of 0.1 atoms / Å3 is shown in black mesh representation.
5 Conclusions
Overall, our findings illuminate the intricate mechanisms underlying the multi-faceted effects of arginine on hydrophobic polymer collapse. Arginine was observed to increase the favorability of many-body hydrophobic interactions, a key factor in protein stabilization. Our observations reveal a nuanced interplay in the impact of arginine on hydrophobic interactions, teetering on the edge of a mechanistic flip. At low concentrations, direct sidechain-driven interactions dominate, shifting to indirect backbone-driven effects at high concentrations.
The simultaneous presence of competing direct and indirect effects implies that changes in the chemistry of a protein surface, the addition of co-additives to solution, or differences in sample preparation may cause significant changes in the mechanism of action of arginine. A shift towards the direct mechanism risks guanidinium-like denaturation of native proteins by disrupting electrostatic and hydrogen-bonding interactions. Conversely, a shift towards the indirect mechanism may yield glycine-like stabilization of native proteins through preferential hydration. Indeed, we illustrated examples of the context-dependent effects of arginine through models of protein (HEWL) and virus (PPV) stability. We observed that arginine destabilizes PPV across a wide concentration range, which we attribute to destabilizing direct arginine-PPV interactions. This mechanism resembles the effects of arginine on a charged polymer model, suggesting that in both cases, arginine-induced destabilization via charge-charge interactions outweighs arginine-induced stabilization of hydrophobic interactions. We further found that arginine destabilizes HEWL at low concentrations and stabilizes HEWL at high concentrations, which we found is associated with an increased exclusion of arginine from the HEWL surface. This mechanism closely resembles the concentration-dependence of the indirect mechanism of arginine in a hydrophobic polymer model. Hence, at high concentrations, HEWL destabilization due to arginine charge-charge interactions may be opposed by an overall stabilization of hydrophobic interactions.
Together, our results suggest that arginine is uniquely situated for use in formulations due to its tunable, context-dependent properties. Hence, while arginine may not be considered the universal stabilizer it once was, its balance between direct- and indirect-driven stabilization of hydrophobic interactions solidifies its significance in formulation design.
Author Contributions
J.Z.: Investigation - Computational, Validation, Formal Analysis, Visualization, Writing - Original Draft, Writing - Review and Editing; P.M.: Investigation - Computational, Validation, Writing - Review and Editing; I.T.: Conceptualization, Investigation - Experimental, Writing - Review and Editing; X.Z.: Investigation - Experimental, Writing - Review and Editing; C.H.: Conceptualization, Funding Acquisition, Resources, Writing - Review and Editing; S.P.: Conceptualization, Funding Acquisition, Resources, Writing - Review and Editing; S.S.: Conceptualization, Methodology, Validation, Supervision, Project Administration, Funding Acquisition, Resources, Writing - Review and Editing
Conflicts of interest
There are no conflicts to declare.
Acknowledgements
This material is based upon work supported by the National Science Foundation under DMREF Grant Nos. 2118788, 2118693, and 2118638.